Find the tension in the horizontal part of the supporting cable, and the compression in the boom, elevated to 25.0° , of the crane that lifts a 8900. N load.

1
Expert's answer
2022-03-06T18:06:32-0500

"T_1=T\cos\alpha=8066~N,"

"T_2=T\sin\alpha=3761~N."